OpenStudy (anonymous):

One number is 7 more than another number. The sum of the numbers is 63. What are the numbers?

OpenStudy (anonymous):

x = y + 7 x + y = 63 Can you solve from here?

OpenStudy (anonymous):

See, Let the one number be y and let the other number be : x So y = x + 7 ---------------1 x + y = 63 So, put the value of y in it: x + x + 7 = 63 2x = 56 x = 28 So, the other number is: y = x + 7 = 28 + 7 = 35 So, the numbers are 28 and 35..